# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ Abstract
In this paper, we investigate the finite-sample statistical rates of distributional temporal difference (TD) learning with linear function approximation. The aim of distributional TD learning is to estimate the return distribution of a discounted Markov decision process for a given policy $\pi$. Prior works on statistical analysis of distributional TD learning mainly focus on the tabular case. In contrast, we first consider the linear function approximation setting and derive sharp finite-sample rates. Our theoretical results demonstrate that the sample complexity of linear distributional TD learning matches that of the classic linear TD learning. This implies that, with linear function approximation, learning the full distribution of the return using streaming data is no more difficult than learning its expectation (i.e. the value function). To derive tight sample complexity bounds, we conduct a fine-grained analysis of the linear-categorical Bellman equation, and employ the exponential stability arguments for products of random matrices. Our findings provide new insights into the statistical efficiency of distributional reinforcement learning algorithms.
@@ 1 Introduction
Distributional policy evaluation (Morimura et al., 2010; Bellemare et al., 2017, 2023), which aims to estimate the return distribution of a policy in a Markov decision process (MDP), is crucial for many uncertainty-aware or risk-sensitive tasks (Lim and Malik, 2022; Kastner et al., 2023). Unlike classic policy evaluation that only focuses on expected returns (i.e. value functions), distributional policy evaluation captures uncertainty and risk by considering the full distributional information. To solve distributional policy evaluation problem, the seminal work of Bellemare et al. (2017) proposed distributional temporal difference (TD) learning, which can be viewed as an extension of classic TD learning (Sutton, 1988).
@@
While classic TD learning was extensively studied (Bertsekas and Tsitsiklis, 1995; Tsitsiklis and Van Roy, 1996; Bhandari et al., 2018; Dalal et al., 2018; Patil et al., 2023; Li et al., 2024a, b; Chen et al., 2024; Samsonov et al., 2024a, b; Wu et al., 2024), the theoretical understanding of distributional TD learning remains relatively underdeveloped. Recent works Rowland et al. (2018); Böck and Heitzinger (2022); Zhang et al. (2023); Rowland et al. (2024a, b); Peng et al. (2024) have analyzed distributional TD learning (or its model-based variants) in the tabular setting. Especially, Rowland et al. (2024b) and Peng et al. (2024) demonstrated that in the tabular setting, statistically speaking, learning the return distribution (in terms of the $1$-Wasserstein distance [^1] Solving distributional policy evaluation $\varepsilon$-accurately in the $1$-Wasserstein distance sense is harder than solving classic policy evaluation $\varepsilon$-accurately, as the absolute difference of value functions is always bounded by the $1$-Wasserstein distance between return distributions. See (Appendix B Peng et al., 2024) for more details.) is as easy as learning its expectation. However, in practical scenarios, where the state space is often extremely large or continuous, function approximation (Bellemare et al., 2017; Dabney et al., 2018b, a; Yang et al., 2019; Freirich et al., 2019; Yue et al., 2020; Nguyen-Tang et al., 2021; Zhou et al., 2021; Luo et al., 2022; Wenliang et al., 2024; Sun et al., 2024) becomes indispensable. This raises a new open question: When function approximation is introduced, does learning the return distribution remain as statistically efficient as learning its expectation?
@@
To answer this question, we consider the simplest form of function approximation, i.e.  linear function approximation, and investigate the finite-sample performance of linear distributional TD learning. In distributional TD, we need to represent the infinite-dimensional return distributions with some finite-dimensional parametrizations to make the algorithm tractable. Prior works (Bellemare et al., 2019; Lyle et al., 2019; Bellemare et al., 2023) have proposed various linear distributional TD learning algorithms under different parameterizations, namely, categorical and quantile parametrizations. In this paper, we consider the categorical parametrization and propose an improved version of categorical linear TD learning algorithm (Linear-CTD). We then analyze the non-asymptotic convergence rate of Linear-CTD. Our analysis reveals that, with Polyak-Ruppert tail averaging (Ruppert, 1988; Polyak and Juditsky, 1992) and constant step size, the sample complexity of Linear-CTD matches that of the classic linear TD learning (Linear-TD) (Li et al., 2024b; Samsonov et al., 2024b). This result confirms that learning the return distribution is statistically no more difficult than learning its expectation when linear function approximation is introduced.
Notations.
@@
In the following parts of the paper, $(x)_{+}:=\max\left\{{x,0}\right\}$ for any $x\in{\mathbb{R}}$. $\lesssim$ (resp. $\gtrsim$) means no larger (resp. smaller) than up to a multiplicative universal constant, and $a\simeq b$ means $a\lesssim b$ and $a\gtrsim b$ hold simultaneously. The asymptotic notation $f(\cdot)=\widetilde{O}\left({g(\cdot)}\right)$ (resp. $\widetilde{\Omega}\left({g(\cdot)}\right)$) means that $f(\cdot)$ is order-wise no larger (resp. smaller) than $g(\cdot)$, ignoring logarithmic factors of polynomials of $(1-\gamma)^{-1},\lambda_{\min}^{-1},\alpha^{-1}$, $\varepsilon^{-1}$, $\delta^{-1}$, $K$, $\left\|{{\bm{\psi}}^{\star}}\right\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}$, $\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$. We will explain the meaning of the notations once we first encounter them. In addition,re the sample complexity more carefully, sometimes we will not ignore the polynomials of specific logarithmic terms in $\widetilde{O}\left({\cdot}\right)$ (resp. $\widetilde{\Omega}\left({\cdot}\right)$), we will emphasize this point when necessary.
@@
We denote by $\otimes$ the Kronecker product (see Appendix A for properties of Kronecker product), $\bm{1}_{K}\in{\mathbb{R}}^{K}$ the all-ones vector, $\bm{0}_{K}\in{\mathbb{R}}^{K}$ the all-zeros vector, ${\bm{I}}_{K}\in{\mathbb{R}}^{K\times K}$ the identity matrix, $\left\|{{\bm{u}}}\right\|$ the Euclidean norm of any vector ${\bm{u}}$, $\left\|{{\bm{B}}}\right\|$ the spectral norm of any matrix ${\bm{B}}$. Let $\left\|{{\bm{u}}}\right\|_{B}:=\sqrt{{\bm{u}}^{\intercal}{\bm{B}}{\bm{u}}}$ for any vector ${\bm{u}}$ and positive semi-definite (PSD) matrix ${\bm{B}}$. $\bm{B}_{1}\preccurlyeq\bm{B}_{2}$ stands for $\bm{B}_{2}-\bm{B}_{1}$ is PSD for any symmetric matrices $\bm{B}_{1},\bm{B}_{2}$. And $\prod_{k=1}^{t}\bm{B}_{k}$ is defined as $\bm{B}_{t}\bm{B}_{t-1}\cdots\bm{B}_{1}$ for any matrices $\left\{{\bm{B}_{k}}\right\}_{k=1}^{t}$. For any matrix ${\bm{B}}=({\bm{b}}(1),\cdots,{\bm{b}}(n))\in{\mathbb{R}}^{m\times n}$, we define its vectorization as $\mathsf{vec}({\bm{B}})=\left({{\bm{b}}(1)^{\intercal},\cdots,{\bm{b}}(n)^{\intercal}}\right)^{\intercal}\in{\mathbb{R}}^{mn}$. Suppose $A$ is some set. We denote by $\Delta(A)$ the set of all probability distributions over $A$. And we denote $\Delta\left({\left[{0,(1-\gamma)^{-1}}\right]}\right)$ by ${\mathscr{P}}$.
1.1 Contributions
Our contribution includes two folds.
@@ 1.1 Contributions
Algorithmically, we propose an improved version of linear categorical TD learning algorithm (Linear-CTD). Instead of using stochastic semi-gradient descent (SSGD) to update the parameter as in (Bellemare et al., 2019; Lyle et al., 2019; Bellemare et al., 2023), we directly translate the linear-categorical projected Bellman equation into a linear system and apply linear stochastic approximation (LSA) to solve it. The resulting Linear-CTD can be viewed as a preconditioned version of the vanilla linear categorical TD learning algorithm proposed in (Section 9.6 Bellemare et al., 2023). By introducing a preconditioner, our Linear-CTD achieve a finite-sample rate independent of the number of supports $K$ in the categorical parameterization, a result the vanilla form cannot attain. We also note that our proposed Linear-CTD is similar to the algorithm in (Lyle et al., 2019), with the distinction that our approach ensures the total mass of the estimated return distributions always be $1$.
@@
Theoretically, we establish the first non-asymptotic guarantees for distributional TD learning with linear function approximation. Specifically, we show that, in the generative model setting, with Polyak-Ruppert tail averaging and instance-dependent constant step size,
@@
$T=\widetilde{O}\left({\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)^{2}\lambda_{\min}}\left({\frac{1}{\varepsilon^{2}}+\frac{1}{\lambda_{\min}}}\right)}\right)$
[…]
@@
online interactions with the environment are sufficient to guarantee Linear-CTD to yield an $\varepsilon$-accurate estimator with high probability, when the error is measured by the $\mu_{\pi}$-weighted $1$-Wasserstein distance. We also provide sample complexity bounds when we take the instance-independent (i.e. not dependent on unknown quantity) step size, and in the Markovian setting. Our sample complexity bounds match those of classic Linear-TD with constant step size (Li et al., 2024b; Samsonov et al., 2024b), confirming the equal statistical tractability of distributional and classic value-based policy evaluation. To establish these theoretical results, we analyze the linear-categorical Bellman equation in details, and apply the exponential stability argument proposed in (Samsonov et al., 2024b). Our analysis of the linear-categorical Bellman equation lays the foundation for subsequent algorithmic and theoretical advances in distributional reinforcement learning with function approximation.
1.2 Related Work
Distributional Reinforcement Learning.
Distributional TD learning was first proposed in (Bellemare et al., 2017). Following the distributional perspective in (Bellemare et al., 2017), Qu et al. (2019) proposed a distributional version of the gradient TD learning algorithm, Tang et al. (2022) proposed a distributional version of multi-step TD learning, Tang et al. (2024) proposed a distributional version of off-policy Q($\lambda$) and TD($\lambda$) algorithms, and Wu et al. (2023) proposed a distributional version of fitted Q evaluation to solve the distributional offline policy evaluation problem. Wiltzer et al. (2024b) proposed an approach for evaluating the return distributions for all policies simultaneously when the reward is deterministic or in the finite-horizon setting. Wiltzer et al. (2024a) studied distributional policy evaluation in the multivariate reward setting and proposed corresponding TD learning algorithms. Beyond the tabular setting, Bellemare et al. (2019); Lyle et al. (2019); Bellemare et al. (2023) proposed various distributional TD learning algorithms with linear function approximation under different parametrizations.
A series of recent studies have focused on the theoretical properties of distributional TD learning. Rowland et al. (2018); Böck and Heitzinger (2022); Zhang et al. (2023); Rowland et al. (2024a, b); Peng et al. (2024) analyzed the asymptotic and non-asymptotic convergence of distributional TD learning (or its model-based variants) in the tabular setting. Among these works, Rowland et al. (2024b); Peng et al. (2024) established that in the tabular setting, learning the full return distribution is statistically as easy as learning its expectation, in the model-based and model-free setting respectively. And Bellemare et al. (2019) provided an asymptotic convergence result for categorical TD learning with linear function approximation.
Beyond the problem of distributional policy evaluation, Rowland et al. (2023); Wang et al. (2023, 2024) showed that, theoretically, classic value-based reinforcement learning could benefit from distributional reinforcement learning. Bäuerle and Ott (2011); Chow and Ghavamzadeh (2014); Marthe et al. (2023); Noorani et al. (2023); Pires et al. (2025) considered optimizing statistical functionals of the return, and proposed algorithms to solve this harder problem.
Stochastic Approximation.
Our Linear-CTD falls into the category of LSA. The classic TD learning, as one of the most classic LSA problems, has been extensively studied (Bertsekas and Tsitsiklis, 1995; Tsitsiklis and Van Roy, 1996; Bhandari et al., 2018; Dalal et al., 2018; Patil et al., 2023; Duan and Wainwright, 2023; Li et al., 2024a, b; Samsonov et al., 2024a; Wu et al., 2024). Among these works, Li et al. (2024b); Samsonov et al. (2024b) provided the tightest bounds for Linear-TD with constant step sizes, which is also considered in our paper. While Wu et al. (2024) established the tightest bounds for Linear-TD with polynomial-decaying step sizes.
For general stochastic approximation problems, extensive works including (Lakshminarayanan and Szepesvari, 2018; Srikant and Ying, 2019; Mou et al., 2020, 2022; Huo et al., 2023; Li et al., 2023b; Durmus et al., 2024; Samsonov et al., 2024b; Chen et al., 2024) have provided solid theoretical understandings.
1.3 Organization
@@
The remainder of this paper is organized as follows. In Section 2, we recap Linear-TD, categorical parametrization of return distributions and categorical TD learning. In Section 3, we introduce the linear-categorical parametrization, and use the linear-categorical projected Bellman equation to derive the Linear-CTD algorithm. In Section 4, we employ the exponential stability arguments to analyze the statistical efficiency of Linear-CTD in both the generative model setting and the Markovian setting. In Section 6, we conclude our work. The details of our proof are given in the appendices.
@@ 2 Backgrounds
2.1 Policy Evaluation and TD learning
@@
A discounted MDP is defined by a 5-tuple $M=\langle{\mathcal{S}},{\mathcal{A}},{\mathcal{P}}_{R},P,\gamma\rangle$. We assume the state space ${\mathcal{S}}$ and action space ${\mathcal{A}}$ are both Polish spaces, namely a complete separable metric space. ${{\mathcal{P}}_{R}}$ is the distribution of rewards, i.e. ${\mathcal{P}}_{R}(\cdot|s,a)\in\Delta\left({[0,1]}\right)$ for any state-action pair $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$ (we assume all rewards is bounded). ${P}$ is the transition dynamics, i.e. $P(\cdot|s,a)\in\Delta\left({{\mathcal{S}}}\right)$ for any $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$. And $\gamma\in(0,1)$ is the discount factor.
@@
Given a policy $\pi\colon{\mathcal{S}}\to\Delta\left({{\mathcal{A}}}\right)$ and an initial state $s_{0}=s\in{\mathcal{S}}$, a random trajectory $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$ can be sampled: $a_{t}\mid s_{t}\sim\pi(\cdot\mid s_{t})$, $r_{t}\mid(s_{t},a_{t})\sim{\mathcal{P}}_{R}({\cdot}\mid s_{t},a_{t})$, ${s_{t+1}}\mid{(s_{t},a_{t})}\sim P({\cdot}\mid{s_{t},a_{t}})$ for any $t\in{\mathbb{N}}$. We assume the Markov chain $\left\{{s_{t}}\right\}_{t=0}^{\infty}$ has a unique stationary distribution $\mu_{\pi}\in\Delta({\mathcal{S}})$. We define the return of the trajectory by $G^{\pi}(s):=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$. The value function $V^{\pi}(s)$ is the expectation of $G^{\pi}(s)$, and ${\bm{V}}^{\pi}:=\left({V^{\pi}(s)}\right)_{s\in{\mathcal{S}}}$. It is known that ${\bm{V}}^{\pi}$ satisfies the Bellman equation:
$V^{\pi}(s)={\mathbb{E}}_{a\sim\pi(\cdot\mid s),r\sim{\mathcal{P}}_{R}({\cdot}\mid s,a),s^{\prime}\sim P(\cdot\mid s,a)}\left[{r+\gamma V^{\pi}(s^{\prime})}\right],\quad\forall s\in{\mathcal{S}},$ (1)
@@
or in a compact form ${\bm{V}}^{\pi}={\bm{T}}^{\pi}{\bm{V}}^{\pi}$, where ${\bm{T}}^{\pi}\colon{\mathbb{R}}^{\mathcal{S}}\to{\mathbb{R}}^{\mathcal{S}}$ is called the Bellman operator.
@@
In the task of policy evaluation, we aim to find ${\bm{V}}^{\pi}$ for some given policy $\pi$. Since ${\bm{V}}^{\pi}$ is the unique solution of the Bellman equation, the problem is reduced to solving the Bellman equation. However, in practical applications ${\bm{T}}^{\pi}$ is usually unknown and the agent only has access to the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$. In this circumstance, we may solve the Bellman equation through LSA. Specifically, at the $t$-th time-step, the updating scheme is
${V}_{t}(s_{t})=V_{t-1}(s_{t})-\alpha\left({V_{t-1}(s_{t})-r_{t}-\gamma V_{t-1}(s_{t+1})}\right).$
@@
We expect ${\bm{V}_{t}}$ will converge to ${\bm{V}}^{\pi}$ as $t$ tends to infinity. This algorithm is known as the TD learning.
2.2 Linear Function Approximation and Linear TD Learning
@@
In this section, we introduce linear function approximation and briefly review Linear-TD. To be concrete, we assume there is a $d$-dimensional feature for each state $s\in{\mathcal{S}}$, which is given by the feature map ${\bm{\phi}}\colon{\mathcal{S}}\to{\mathbb{R}}^{d}$. We consider the linear function approximation of value functions:
${\mathscr{V}}_{{\bm{\phi}}}:=\left\{{\bm{V}}_{{\bm{\psi}}}=\left({V_{{\bm{\psi}}}(s)}\right)_{s\in{\mathcal{S}}}\colon V_{{\bm{\psi}}}(s)={\bm{\psi}}^{\intercal}{\bm{\phi}}(s),{\bm{\psi}}\in{\mathbb{R}}^{d}\right\}\subset{\mathbb{R}}^{\mathcal{S}},$

### v2

# A Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

@@ Abstract
In this paper, we study the finite-sample statistical rates of distributional temporal difference (TD) learning with linear function approximation. The aim of distributional TD learning is to estimate the return distribution of a discounted Markov decision process for a given policy $\pi$. Previous works on statistical analysis of distributional TD learning mainly focus on the tabular case. In contrast, we first consider the linear function approximation setting and derive sharp finite-sample rates. Our theoretical results demonstrate that the sample complexity of linear distributional TD learning matches that of classic linear TD learning. This implies that, with linear function approximation, learning the full distribution of the return from streaming data is no more difficult than learning its expectation (value function). To derive tight sample complexity bounds, we conduct a fine-grained analysis of the linear-categorical Bellman equation and employ the exponential stability arguments for products of random matrices. Our results provide new insights into the statistical efficiency of distributional reinforcement learning algorithms.
@@ 1 Introduction
Distributional policy evaluation [36, 2, 4], which aims to estimate the return distribution of a policy in an Markov decision process (MDP), is crucial for many uncertainty-aware or risk-sensitive tasks [30, 21]. Unlike the classic policy evaluation that only focuses on expected returns (value functions), distributional policy evaluation captures uncertainty and risk by considering the full distributional information. To solve a distributional policy evaluation problem, in the seminal work [2] proposed distributional temporal difference (TD) learning, which can be viewed as an extension of classic TD learning [58].
@@
Although classic TD learning has been extensively studied [5, 61, 6, 14, 41, 25, 26, 10, 53, 54, 69], the theoretical understanding of distributional TD learning remains relatively underdeveloped. Recent works [48, 7, 73, 50, 51, 43] have analyzed distributional TD learning (or its model-based variants) in the tabular setting. Especially, Rowland et al. [51] and Peng et al. [43] demonstrated that in the tabular setting, learning the return distribution (in terms of the $1$-Wasserstein distance [^1] Solving distributional policy evaluation $\varepsilon$-accurately in the $1$-Wasserstein distance sense is harder than solving classic policy evaluation $\varepsilon$-accurately, as the absolute difference of value functions is always bounded by the $1$-Wasserstein distance between return distributions. ) is statistically as easy as learning its expectation. However, in practical scenarios, where the state space is extremely large or continuous, the function approximation [13, 12, 70, 17, 71, 39, 74, 31, 65, 57] becomes indispensable. This raises a new open question: When function approximation is employed, does learning the return distribution remain as statistically efficient as learning its expectation?
@@
To answer this question, we consider the simplest form of function approximation, i.e., linear function approximation, and investigate the finite-sample performance of linear distributional TD learning. In distributional TD learning, we need to represent the infinite-dimensional return distributions with some finite-dimensional parametrizations to make the algorithm tractable. Previous works [3, 32, 4] have proposed various linear distributional TD learning algorithms under different parameterizations, namely categorical and quantile parametrizations. In this paper, we consider the categorical parametrization and propose an improved version of the linear-categorical TD learning algorithm (Linear-CTD). We then analyze the non-asymptotic convergence rate of Linear-CTD. Our analysis reveals that, with the Polyak-Ruppert tail averaging [52, 45] and a proper constant step size, the sample complexity of Linear-CTD matches that of classic linear TD learning (Linear-TD) [26, 54]. Thus, this confirms that learning the return distribution is statistically no more difficult than learning its expectation when the linear function approximation is employed.
Notation.
@@
In the following parts of the paper, $(x)_{+}:=\max\left\{{x,0}\right\}$ for any $x\in{\mathbb{R}}$. $\lesssim$” (resp. $\gtrsim$”) means no larger (resp. smaller) than up to a multiplicative universal constant, and $a\simeq b$ means $a\lesssim b$ and $a\gtrsim b$ hold simultaneously. The asymptotic notation $f(\cdot)=\widetilde{{\mathcal{O}}}\left({g(\cdot)}\right)$ (resp. $\widetilde{\Omega}\left({g(\cdot)}\right)$) means that $f(\cdot)$ is order-wise no larger (resp. smaller) than $g(\cdot)$, ignoring logarithmic factors of polynomials of $(1-\gamma)^{-1},\lambda_{\min}^{-1},\alpha^{-1}$, $\varepsilon^{-1}$, $\delta^{-1}$, $K$, $\left\|{{\bm{\psi}}^{\star}}\right\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}$, $\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$. We will explain the concrete meaning of the notation once we have encountered them for the first time.
@@
We denote by $\delta_{x}$ the Dirac measure at $x\in{\mathbb{R}}$, $\mathds{1}$ the indicator function, $\otimes$ the Kronecker product (see Appendix A), $\bm{1}_{K}{\in}{\mathbb{R}}^{K}$ the all-ones vector, $\bm{0}_{K}{\in}{\mathbb{R}}^{K}$ the all-zeros vector, ${\bm{I}}_{K}{\in}{\mathbb{R}}^{K{\times}K}$ the identity matrix, $\left\|{{\bm{u}}}\right\|$ the Euclidean norm of any vector ${\bm{u}}$, $\left\|{{\bm{B}}}\right\|$ the spectral norm of any matrix ${\bm{B}}$, and $\left\|{{\bm{u}}}\right\|_{B}:=\sqrt{{\bm{u}}^{\top}{\bm{B}}{\bm{u}}}$ when ${\bm{B}}$ is positive semi-definite (PSD). $\bm{B}_{1}{\preccurlyeq}\bm{B}_{2}$ stands for $\bm{B}_{2}{-}\bm{B}_{1}$ is PSD for any symmetric matrices $\bm{B}_{1},\bm{B}_{2}$. And $\prod_{k=1}^{t}\bm{B}_{k}$ is defined as $\bm{B}_{t}\bm{B}_{t{-}1}\cdots\bm{B}_{1}$ for any matrices $\left\{{\bm{B}_{k}}\right\}_{k=1}^{t}$ with appropriate sizes. For any matrix ${\bm{B}}{=}[{\bm{b}}(1),\ldots,{\bm{b}}(n)]{\in}{\mathbb{R}}^{m{\times}n}$, we define its vectorization as $\mathsf{vec}({\bm{B}}){=}({\bm{b}}(1)^{\top},\ldots,{\bm{b}}(n)^{\top})^{\top}{\in}{\mathbb{R}}^{mn}$. Given a set $A$, we denote by $\Delta(A)$ the set of all probability distributions over $A$. For simplicity, we abbreviate $\Delta([0,(1{-}\gamma)^{-1}])$ as ${\mathscr{P}}$.
Contributions.
@@
Our contribution is two-fold: in algorithms and in theory. Algorithmically, we propose an improved version of the linear-categorical TD learning algorithm (Linear-CTD). Rather than using stochastic semi-gradient descent to update the parameter as in Bellemare et al. [3], Lyle et al. [32], Bellemare et al. [4], we directly formulate the linear-categorical projected Bellman equation into a linear system and apply a linear stochastic approximation to solve it. The resulting Linear-CTD can be viewed as a preconditioned version [9, 29] of the vanilla linear categorical TD learning algorithm proposed in Bellemare et al. [4, Section 9.6]. By introducing a preconditioner, our Linear-CTD achieves a finite-sample rate independent of the number of supports $K$ in the categorical parameterization, which the vanilla version cannot attain. We provide both theoretical and experimental evidence to demonstrate this advantage of our Linear-CTD.
@@
Theoretically, we establish the first non-asymptotic guarantees for distributional TD learning with the linear function approximation. Specifically, we show that in the generative model setting, with the Polyak-Ruppert tail averaging and a constant step size, we need
@@
$T=\widetilde{{\mathcal{O}}}\left({\left({{\varepsilon^{-2}}+{\lambda_{\min}^{-1}}}\right)(1-\gamma)^{-2}\lambda_{\min}^{-1}\left({{K^{-1}(1-\gamma)^{-2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}\right)}\right)$
@@
online interactions with the environment to ensure Linear-CTD yields a $\varepsilon$-accurate estimator with high probability, when the error is measured by the $\mu_{\pi}$-weighted $1$-Wasserstein distance. We also extend the result to the Markovian setting. Our sample complexity bounds match those of the classic Linear-TD with a constant step size [26, 54], confirming the same statistical tractability of distributional and classic value-based policy evaluations. To establish these theoretical results, we analyze the linear-categorical Bellman equation in detail and apply the exponential stability argument proposed in Samsonov et al. [54]. Our analysis of the linear-categorical Bellman equation lays the foundation for subsequent algorithmic and theoretical advances in distributional reinforcement learning with function approximation.
Organization.
@@
The remainder of this paper is organized as follows. In Section 2, we recap Linear-TD and tabular categorical TD learning. In Section 3, we introduce the linear-categorical parametrization, and use the linear-categorical projected Bellman equation to derive Linear-CTD. In Section 4, we employ the exponential stability arguments to analyze the statistical efficiency of Linear-CTD. The proof is outlined in Section 5. In Section 6, we conclude our work. See Appendix B for more related work. In Appendix F, we empirically validate the convergence of Linear-CTD and compare it with prior algorithms through numerical experiments, confirming our theoretical findings. Details of the proof are given in the appendices.
@@ 2 Backgrounds
In this section, we recap the basics of policy evaluation and distributional policy evaluation tasks.
2.1 Policy Evaluation
@@
A discounted MDP is defined by a $4$-tuple $M=\langle{\mathcal{S}},{\mathcal{A}},{\mathcal{P}},\gamma\rangle$. We assume that the state space ${\mathcal{S}}$ and the action space ${\mathcal{A}}$ are both Polish spaces, namely complete separable metric spaces. ${{\mathcal{P}}(\cdot,\cdot\mid s,a)}$ is the joint distribution of reward and next state condition on $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$. We assume that all rewards are bounded random variables in $[0,1]$. And $\gamma\in(0,1)$ is the discount factor.
@@
Given a policy $\pi\colon{\mathcal{S}}\to\Delta\left({{\mathcal{A}}}\right)$ and an initial state $s_{0}=s\in{\mathcal{S}}$, a random trajectory $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$ can be sampled: $a_{t}\mid s_{t}\sim\pi(\cdot\mid s_{t})$, $(r_{t},s_{t+1})\mid(s_{t},a_{t})\sim{\mathcal{P}}({\cdot,\cdot}\mid s_{t},a_{t})$, for any $t\in{\mathbb{N}}$. We assume the Markov chain $\left\{{s_{t}}\right\}_{t=0}^{\infty}$ has a unique stationary distribution $\mu_{\pi}\in\Delta({\mathcal{S}})$. We define the return of the trajectory as $G^{\pi}(s):=\sum_{t=0}^{\infty}\gamma^{t}r_{t}$. The value function $V^{\pi}(s)$ is the expectation of $G^{\pi}(s)$, and ${\bm{V}}^{\pi}:=\left({V^{\pi}(s)}\right)_{s\in{\mathcal{S}}}{\in}{\mathbb{R}}^{\mathcal{S}}$. It is known that ${\bm{V}}^{\pi}$ satisfies the Bellman equation:
$V^{\pi}(s)={\mathbb{E}}_{a\sim\pi(\cdot\mid s),(r,s^{\prime})\sim{\mathcal{P}}(\cdot,\cdot\mid s,a)}[r+\gamma V^{\pi}(s^{\prime})],\quad\forall s\in{\mathcal{S}},$ (1)
@@
or in a compact form ${\bm{V}}^{\pi}={\bm{T}}^{\pi}{\bm{V}}^{\pi}$, where ${\bm{T}}^{\pi}\colon{\mathbb{R}}^{\mathcal{S}}\to{\mathbb{R}}^{\mathcal{S}}$ is called the Bellman operator. In the task of policy evaluation, we aim to find the unique solution ${\bm{V}}^{\pi}$ of the equation for some given policy $\pi$.
Tabular TD Learning.
@@
The policy evaluation problem is reduced to solving the Bellman equation. However, in practical applications ${\bm{T}}^{\pi}$ is usually unknown and the agent only has access to the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$. In this circumstance, we can solve the Bellman equation through linear stochastic approximation (LSA). Specifically, in the $t$-th time-step the updating scheme is
${V}_{t}(s_{t})\leftarrow V_{t-1}(s_{t})-\alpha\left({V_{t-1}(s_{t})-r_{t}-\gamma V_{t-1}(s_{t+1})}\right),\quad V_{t}(s)\leftarrow V_{t-1}(s),\ \forall s\neq s_{t}.$ (2)
@@
We expect ${\bm{V}_{t}}$ to converge to ${\bm{V}}^{\pi}$ as $t$ tends to infinity. This algorithm is known as TD learning, however, it is computationally tractable only in the tabular setting.
Linear Function Approximation and Linear TD Learning.
@@
In this part, we introduce linear function approximation and briefly review the more practical Linear-TD. To be concrete, we assume there is a $d$-dimensional feature vector for each state $s\in{\mathcal{S}}$, which is given by the feature map ${\bm{\phi}}\colon{\mathcal{S}}\to{\mathbb{R}}^{d}$. We consider the linear function approximation of value functions:
${\mathscr{V}}_{{\bm{\phi}}}:=\left\{{\bm{V}}_{{\bm{\psi}}}=\left({V_{{\bm{\psi}}}(s)}\right)_{s\in{\mathcal{S}}}\colon V_{{\bm{\psi}}}(s)={\bm{\phi}}(s)^{\top}{\bm{\psi}},{\bm{\psi}}\in{\mathbb{R}}^{d}\right\}\subset{\mathbb{R}}^{\mathcal{S}},$ (3)